OpenStudy (anonymous):

x approaching positive infinity (2-y)/(sqrt(7+6y^2))

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do this in your head

OpenStudy (misty1212):

hmm i assume it is \(y\to \infty\) right?

OpenStudy (anonymous):

ha yes. sorry about that

OpenStudy (misty1212):

ignore all but the highest power, think of this as \[\frac{-x}{\sqrt{6x^2}}=\frac{-x}{\sqrt6 x}\]

OpenStudy (misty1212):

cancel and get \[\frac{-1}{\sqrt6}\]

OpenStudy (anonymous):

wow. Thank you so much for breaking that down. You're a wiz

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]